An ideal gas at 45 psig and 80°F is heated in a closed container to 130°F. What is the final pressure?   


1
Expert's answer
2021-10-15T10:24:15-0400

In a closed container volume is constant. At constant volume by Charle’s law

"\frac{p_2}{p_1} = \frac{T_2}{T_1} \\

T_1 = 80 \;°F = 299.8 \;K \\

T_2 = 130 \;°F = 327.5 \;K \\

p_2 = \frac{T_2p_1}{T_1} \\

p_2= \frac{327.5 \times 45}{299.8} = 49.15 \;psi"